# Corresponding Authors

### v1

@@ 6 Conclusion
In this paper, we proposed MoMo, a disentangled motion modeling framework for perceptual video frame interpolation. Our approach mainly focuses on modeling the intermediate motions between frames, with explicit supervision on the motions. We introduced motion diffusion model, which generates intermediate bi-directional flows necessary to synthesize the target frame. We also presented a novel architecture for diffusion models, tailored for optical flow generation, which greatly improves both performance and computational efficiency. Extensive experiments confirm that our method achieve state-of-the-art quality on multiple benchmarks.
6.0.1 Acknowledgement
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A3B1077720) and the BK21 FOUR program of the Education and the Research Program for Future ICT Pioneers, Seoul National University in 2024.
Appendix 0.A Implementation Details
0.A.1 Perception-oriented loss
As mentioned in Sec. 3.2, we specify the objective function we use for stage 1 training. Along with the $L_{1}$ pixel reconstruction loss, we use an LPIPS-based perceptual loss, which computes the $L_{2}$ distance in the deep feature space of AlexNet [29]. This loss is well-known for high correlation with human judgements. Next, we exploit the style loss [15] $\mathcal{L}_{G}$ as its effectiveness has been proved in a recent work [52]. This loss computes the $L_{2}$ distance of feature correlations extracted from the VGG-19 network [62]:
@@ 0.A.1 Perception-oriented loss
Here, $\alpha_{n}$ denotes the weighting hyper-parameter of the $n$-th selected layer. Denoting the feature map of frame $I_{\tau}$ extracted from $n$-th selected layer of the VGG [62] network as $\phi_{n}(I_{\tau})\in\mathbb{R}^{H\times W\times C}$, the Gram matrix of frame $I_{\tau}$ at the $n$-th feature space, $G_{n}(I_{\tau})\in\mathbb{R}^{C\times C}$ can be acquired as follows:
@@
0.A.2 Training Details
We elaborate on our training parameters as mentioned in Sec 4.1.
@@ 0.A.2 Training Details
We employ the AdamW optimizer [34], setting the weight decay to $10^{-4}$ and the batch size to 32 both in the training process of $\mathcal{G}$ and $\mathcal{F}$. We train the synthesis model $\mathcal{G}$ for a total of 200 epochs, with a fixed learning rate of $2\times 10^{-4}$. For the first 150 epochs, we set the hyper-parameters to $\lambda_{1}=1,\lambda_{p}=0,\lambda_{G}=0$. After that, we use $\lambda_{1}=1,\lambda_{p}=1,\lambda_{G}=20$ for the last 50 epochs. Once the synthesis model is fully trained, we fine-tune the teacher flow model $\mathcal{F}$ for 100 epochs, with its learning rate fixed to $10^{-4}$. We set hyper-parameters to $\lambda_{1}=1,\lambda_{p}=1,\lambda_{G}=20$. Both $\mathcal{G}$ and $\mathcal{F}$ benefit from an exponential moving average (EMA) with a 0.999 decay rate. We set the number of pyramids $L=3$ during training, and use $L=\lceil\log_{2}(R/32)\rceil$ for resolution $R$ at inference.
@@
We train our diffusion model for 500 epochs using the AdamW optimizer with a constant learning rate of $2\times 10^{-4}$, weight decay of $10^{-8}$, and batch size of 64, applying an EMA with a 0.9999 decay rate. Given that diffusion models typically operate with data values between $[-1,1]$ but optical flows often exceed this range, we normalize flow values by dividing them by 128. This adjustment ensures flow values to be compatible with the diffusion model’s expected data range, effectively aligning flow values with those of the RGB space, which are similarly normalized. We utilize a linear noise schedule [18] and perform 8 denoising steps using the ancestral DDPM sampler [18] for efficient sampling.
Appendix 0.B Additional Results
0.B.1 Deeper Analysis on Number of Denoising Steps
We experiment on the effect of different number of denoising steps for motion modeling, on the ‘hard’ subset of SNU-FILM.(Tab.6) Although the performance does improve up to 8 steps, the increase is relatively marginal compared to the results on the ‘extreme’ subset. We speculate the reason for this result is due to the smaller ill-posedness of the ‘hard’ subset, which limits the diversity of feasible flows. We claim that the use of more steps and the design choice of diffusion models for motion modeling gets more advantageous as the ill-posedness of motions gets larger.
@@ 0.B.1 Deeper Analysis on Number of Denoising Steps
Table 6: Experiment on the number of denoising steps at inference time. Our experiments show that about 8 steps is enough, and use of more steps exceeding this does not lead to a notable improvement considering the runtime tradeoff.
0.B.2 Full Quantitative Results
We report the full quantitative results including the fidelity metrics such as PSNR and SSIM on the SNU-FILM (Tab. 78), Middlebury (Tab. 9) and Xiph benchmarks (Tab. 10).
Method SNU-FILM-easy SNU-FILM-medium
@@ 0.B.2 Full Quantitative Results
XVFI [61] 38.57 0.9824 0.0251 0.0219 34.56 0.9635 0.0389 0.0304
XVFIv [61] 39.78 0.9865 0.0175 0.0181 35.36 0.9692 0.0322 0.0276
RIFE [22] 40.06 0.9907 0.0181 0.0195 35.75 0.9789 0.0317 0.0289
IFRNet [28] 40.03 0.9905 0.0201 0.0206 35.94 0.9793 0.0320 0.0281
FILM-$\mathcal{L}_{1}$ [52] 39.74 0.9902 0.0184 0.0217 35.81 0.9789 0.0315 0.0316
ABME [49] 39.59 0.9901 0.0222 0.0229 35.77 0.9789 0.0372 0.0344
IFRNet-Large [28] 40.10 0.9906 0.0203 0.0211 36.12 0.9797 0.0321 0.0288
UPRNet [24] 40.37 0.9910 0.0179 0.0200 36.16 0.9797 0.0322 0.0311
AMT-G [32] 38.47 0.9880 0.0325 0.0312 35.39 0.9779 0.0447 0.0395
EMA-VFI [72] 39.52 0.9903 0.0186 0.0204 35.83 0.9795 0.0325 0.0318
UPRNet-LARGE [24] 40.44 0.9911 0.0182 0.0203 36.29 0.9801 0.0334 0.0327
CAIN [6] 39.89 0.9900 0.0197 0.0229 35.61 0.9776 0.0375 0.0347
FILM-$\mathcal{L}_{vgg}$ [52] 39.79 0.9900 0.0123 0.0128 35.77 0.9782 0.0219 0.0183
FILM-$\mathcal{L}_{s}$ [52] 39.68 0.9900 0.0120 0.0124 35.70 0.9781 0.0213 0.0177
LDMVFI [8] 38.68 0.9834 0.0145 0.0130 33.90 0.9703 0.0284 0.0219
@@
[…]
Method SNU-FILM-hard SNU-FILM-extreme
@@
XVFI [61] 29.56 0.9028 0.0668 0.0431 24.89 0.8095 0.1189 0.0638
XVFIv [61] 29.91 0.9073 0.0629 0.0414 24.67 0.8092 0.1257 0.0673
RIFE [22] 30.10 0.9330 0.0657 0.0443 24.84 0.8534 0.1390 0.0764
IFRNet [28] 30.41 0.9358 0.0573 0.0398 25.05 0.8587 0.1176 0.0653
FILM-$\mathcal{L}_{1}$ [52] 30.42 0.9353 0.0568 0.0441 25.17 0.8593 0.1060 0.0632
ABME [49] 30.58 0.9364 0.0658 0.0496 25.42 0.8639 0.1258 0.0747
IFRNet-Large [28] 30.63 0.9368 0.0562 0.0403 25.27 0.8609 0.1131 0.0638
UPRNet [24] 30.67 0.9365 0.0604 0.0452 25.49 0.8627 0.1115 0.0653
AMT-G [32] 30.70 0.9381 0.0680 0.0506 25.64 0.8658 0.1128 0.0686
EMA-VFI [72] 30.79 0.9386 0.0579 0.0457 25.59 0.8648 0.1099 0.0671
UPRNet-LARGE [24] 30.86 0.9377 0.0612 0.0475 25.63 0.8641 0.1109 0.0672
CAIN [6] 29.90 0.9292 0.0885 0.0606 24.78 0.8507 0.1790 0.1042
FILM-$\mathcal{L}_{vgg}$ [52] 30.34 0.9332 0.0443 0.0282 25.11 0.8557 0.0917 0.0471
FILM-$\mathcal{L}_{s}$ [52] 30.29 0.9329 0.0429 0.0268 25.07 0.8550 0.0889 0.0448
LDMVFI [8] 28.51 0.9173 0.0602 0.0379 23.92 0.8372 0.1226 0.0651
@@
[…]
Method Middlebury
PSNR SSIM LPIPS DISTS
XVFI [61] 33.43 0.9557 0.0520 0.0557
XVFIv [61] 36.72 0.9826 0.0169 0.0244
RIFE [22] 37.16 0.9853 0.0162 0.0228
FILM-$\mathcal{L}_{1}$ [52] 37.37 0.9838 0.0173 0.0246
ABME [49] 37.05 0.9845 0.0290 0.0325
IFRNet-Large [28] 36.27 0.9816 0.0285 0.0366
AMT-G [32] 34.23 0.9708 0.0486 0.0533
EMA-VFI [72] 38.32 0.9871 0.0151 0.0218
UPRNet-LARGE [24] 38.09 0.9861 0.0150 0.0209
CAIN [6] 35.11 0.9761 0.0254 0.0383
FILM-${\mathcal{L}_{vgg}}^{{\dagger}}$ [52] 37.28 0.9843 0.0096 0.0148
FILM-$\mathcal{L}_{s}$ [52] 37.38 0.9844 0.0093 0.0140
LDMVFI [8] 34.03 0.9648 0.0195 0.0261
MoMo (Ours) 36.77 0.9806 0.0094 0.0126
[…]
Method Xiph-2K Xiph-4K
@@
XVFI [61] 34.75 0.9560 0.0710 0.0360 32.83 0.9231 0.1222 0.0610
XVFIv [61] 35.17 0.9625 0.0844 0.0418 32.45 0.9274 0.1835 0.0779
RIFE [22] 36.06 0.9642 0.0918 0.0481 33.21 0.9413 0.2072 0.0915
FILM-$\mathcal{L}_{1}$ [52] 36.53 0.9663 0.0906 0.0510 33.83 0.9439 0.1841 0.0884
ABME [49] 36.50 0.9668 0.1071 0.0581 33.72 0.9452 0.2361 0.1108
IFRNet-Large [28] 36.40 0.9646 0.0681 0.0372 33.71 0.9425 0.1364 0.0665
AMT-G [32] 36.29 0.9647 0.1061 0.0563 34.55 0.9472 0.2054 0.1005
EMA-VFI [72] 36.74 0.9675 0.1024 0.0550 34.55 0.9486 0.2258 0.1049
UPRNet-LARGE [24] 37.13 0.9691 0.1010 0.0553 34.57 0.9388 0.2150 0.1017
CAIN [6] 35.18 0.9625 0.1025 0.0533 32.55 0.9398 0.2229 0.0980
FILM-${\mathcal{L}_{vgg}}^{{\dagger}}$ [52] 36.29 0.9626 0.0355 0.0238 33.44 0.9356 0.0754 0.0406
FILM-$\mathcal{L}_{s}$ [52] 36.30 0.9616 0.0330 0.0237 33.37 0.9323 0.0703 0.0385
LDMVFI [8] 33.82 0.9494 0.0420 0.0163 31.39 0.9214 0.0859 0.0359
@@
[…]

### v2

@@ 5 Conclusion
In this paper, we proposed MoMo, a disentangled motion modeling framework for perceptual video frame interpolation. Our approach mainly focuses on modeling the intermediate motions between frames, with explicit supervision on the motions only. We introduced motion diffusion model, which generates intermediate bi-directional flows necessary to synthesize the target frame with a novel architecture tailored for optical flow generation, which greatly improves both performance and computational efficiency. Extensive experiments confirm that our method achieve state-of-the-art quality on multiple benchmarks.
Acknowledgement
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A3B1077720), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(No. 2022R1A5A7083908), and the BK21 FOUR program of the Education and the Research Program for Future ICT Pioneers, Seoul National University in 2024.
Appendix A Implementation Details
A.1 Recurrent Synthesis
We build our synthesis network $\mathcal{S}$ to be of recurrent structure, motivated by the recent trend in video frame interpolation (Sim, Oh, and Kim 2021; Jin et al. 2023; Reda et al. 2022), due to its great efficiency. Let the number of recurrent process be $L-1$, and $\mathcal{S}$ can be expressed as recurrent application of a synthesis process $\mathcal{P}$:
$\mathcal{S}(I_{in},F_{\tau})=\mathcal{P}^{0}(\cdots\mathcal{P}^{L-1}(\hat{I}^{L}_{\tau},I^{L-1}_{in},F^{L-1}_{\tau})\cdots,I^{0}_{in},F^{0}_{\tau}),$ (9)
Next, we take frame $\hat{I}^{l+1}_{\tau}$ and use bicubic upsampling to match the size of level $l$, denoted as $\hat{I}^{l+1\rightarrow l}_{\tau}$. The upsampled frame $\hat{I}^{l+1\rightarrow l}_{\tau}$, along with the two warped frames $I^{l}_{\tau\leftarrow 0},I^{l}_{\tau\leftarrow 1}$ and their corresponding flow maps $F^{l}_{\tau\rightarrow 0},F^{l}_{\tau\rightarrow 1}$ are given to the synthesis module $\mathcal{G}$, which outputs a 4 channel output — 1 channel occlusion mask $M_{0}$ to blend $I_{\tau\leftarrow 0}$ and $I_{\tau\leftarrow 1}$, and 3 channel residual RGB values $\Delta\hat{I_{\tau}}$:
$M^{l}_{0},\Delta\hat{I}^{l}_{\tau}=\mathcal{G}(\hat{I}^{l+1\rightarrow l}_{\tau},I^{l}_{\tau\leftarrow 0},I^{l}_{\tau\leftarrow 1},F^{l}_{\tau\rightarrow 0},F^{l}_{\tau\rightarrow 1}).$ (11)
$\mathcal{P}^{l}(\hat{I}^{l+1}_{\tau},I_{in}^{l},F_{\tau}^{l})=I^{l}_{\tau\leftarrow 0}\odot M^{l}_{0}+I^{l}_{\tau\leftarrow 1}\odot(1-M^{l}_{0})+\Delta\hat{I}^{l}_{\tau}.$ (12)
Note that $\hat{I}^{L}_{\tau}$ is not available for level $\mathcal{P}^{L-1}$, since it is of the highest level. Therefore we equally blend the two warped frames at level $l=L-1$ as a starting point: $\hat{I}^{L\rightarrow L-1}_{\tau}=I^{L-1}_{\tau\leftarrow 0}\odot 0.5+I^{L-1}_{\tau\leftarrow 1}\odot 0.5$.
A.2 Perception-oriented Loss
As mentioned in Sec. 3.2, we specify the objective function we use for stage 1 training. Along with the $L_{1}$ pixel reconstruction loss, we use an LPIPS-based perceptual loss, which computes the $L_{2}$ distance in the deep feature space of AlexNet (Krizhevsky, Sutskever, and Hinton 2012). This loss is well-known for high correlation with human judgements. Next, we exploit the style loss (Gatys, Ecker, and Bethge 2016) $\mathcal{L}_{G}$ as its effectiveness has been proved in a recent work (Reda et al. 2022). This loss computes the $L_{2}$ distance of feature correlations extracted from the VGG-19 network (Simonyan and Zisserman 2014):
@@ A.2 Perception-oriented Loss
Here, $\alpha_{n}$ denotes the weighting hyper-parameter of the $n$-th selected layer. Denoting the feature map of frame $I_{\tau}$ extracted from $n$-th selected layer of the VGG (Simonyan and Zisserman 2014) network as $\phi_{n}(I_{\tau})\in\mathbb{R}^{H\times W\times C}$, the Gram matrix of frame $I_{\tau}$ at the $n$-th feature space, $G_{n}(I_{\tau})\in\mathbb{R}^{C\times C}$ can be acquired as follows:
@@
A.3 Training Details
We elaborate on our training details mentioned in Sec. 4.1.
@@ A.3 Training Details
We employ the AdamW optimizer (Loshchilov and Hutter 2017), setting the weight decay to $10^{-4}$ and the batch size to 32 both in the training process of $\mathcal{G}$ and $\mathcal{F}$. We train the synthesis model $\mathcal{G}$ for a total of 200 epochs, with a fixed learning rate of $2\times 10^{-4}$. For the first 150 epochs, we set the hyper-parameters to $\lambda_{1}=1,\lambda_{p}=0,\lambda_{G}=0$. After that, we use $\lambda_{1}=1,\lambda_{p}=1,\lambda_{G}=20$ for the last 50 epochs. Once the synthesis model is fully trained, we fine-tune the teacher flow model $\mathcal{F}$ for 100 epochs, with its learning rate fixed to $10^{-4}$. We set hyper-parameters to $\lambda_{1}=1,\lambda_{p}=1,\lambda_{G}=20$. Both $\mathcal{G}$ and $\mathcal{F}$ benefit from an exponential moving average (EMA) with a 0.999 decay rate. We set the number of pyramids $L=3$ during training, and use $L=\lceil\log_{2}(R/32)\rceil$ for resolution $R$ at inference.
@@
We train our diffusion model for 500 epochs using the AdamW optimizer with a constant learning rate of $2\times 10^{-4}$, weight decay of $10^{-8}$, and batch size of 64, applying an EMA with a 0.9999 decay rate. Given that diffusion models typically operate with data values between $[-1,1]$ but optical flows often exceed this range, we normalize flow values by dividing them by 128. This adjustment ensures flow values to be compatible with the diffusion model’s expected data range, effectively aligning flow values with those of the RGB space, which are similarly normalized. We utilize a linear noise schedule (Ho, Jain, and Abbeel 2020) and perform 8 denoising steps using the ancestral DDPM sampler (Ho, Jain, and Abbeel 2020) for efficient sampling.
A.4 Inference
Since our motion diffusion model is trained on a well-curated data ranging within 256 resolution, it could suffer from a performance drop when it comes to high resolution videos of large motions which goes beyond the distribution of the training data. To handle these cases, we generate flows at the training resolution by resizing the inputs, followed by post-processing of bicubic upsampling at inference time.
A.5 Input Downsampling
We provide illustrated description of input downsampling (Sec. 3.3) in Fig. A1.
Figure A1: Visualized description of input downsampling in our motion diffusion model.
A.6 Convex Upsampling
We provide a illustrated description of convex upsampling (Sec. 3.3) in Fig. A2.
Figure A2: Visualization of convex upsampling layer. The dimensions of flows are omitted for simplicity. The upsampling weights are predicted for both directions and applied to the bi-directional flows in the same manner.
Appendix B Additional Experiments
B.1 Further Analysis on Denoising Steps
We experiment on the effect of different number of denoising steps for motion modeling, on the ‘hard’ subset of SNU-FILM.(Tab.A1) Although the performance does improve up to 8 steps, the increase is relatively marginal compared to the results on the ‘extreme’ subset. We speculate the reason for this result is due to the smaller ill-posedness of the ‘hard’ subset, which limits the diversity of feasible flows. We claim that the use of more steps and the design choice of diffusion models for motion modeling is more advantageous as the ill-posedness of motions gets larger.
@@ B.1 Further Analysis on Denoising Steps
Table A1: Experiment on the number of denoising steps at inference time. Our experiments show that about 8 steps is enough, and use of more steps exceeding this does not lead to a notable improvement considering the runtime tradeoff.
B.2 Full Quantitative Results
We report the full quantitative results including the fidelity metrics such as PSNR and SSIM on the SNU-FILM (Tab. A2A3), Middlebury, Vimeo90k (Tab. A4) and Xiph benchmarks (Tab. A5).
In addition to the fidelity metrics, we also include the results of a lighter version of our model with 10M parameters, denoted as MoMo-10M.
Method Perception- oriented loss SNU-FILM-easy SNU-FILM-medium
@@ B.2 Full Quantitative Results
ABME (Park, Lee, and Kim 2021) 39.59 0.9901 0.0222 0.0229 35.77 0.9789 0.0372 0.0344
XVFIv (Sim, Oh, and Kim 2021) 39.78 0.9865 0.0175 0.0181 35.36 0.9692 0.0322 0.0276
IFRNet-Large (Kong et al. 2022) 40.10 0.9906 0.0203 0.0211 36.12 0.9797 0.0321 0.0288
RIFE (Huang et al. 2022b) 40.06 0.9907 0.0181 0.0195 35.75 0.9789 0.0317 0.0289
FILM-$\mathcal{L}_{1}$ (Reda et al. 2022) 39.74 0.9902 0.0184 0.0217 35.81 0.9789 0.0315 0.0316
AMT-G (Li et al. 2023) 38.47 0.9880 0.0325 0.0312 35.39 0.9779 0.0447 0.0395
EMA-VFI (Zhang et al. 2023) 39.52 0.9903 0.0186 0.0204 35.83 0.9795 0.0325 0.0318
UPRNet-LARGE (Jin et al. 2023) 40.44 0.9911 0.0182 0.0203 36.29 0.9801 0.0334 0.0327
CAIN (Choi et al. 2020) 39.89 0.9900 0.0197 0.0229 35.61 0.9776 0.0375 0.0347
FILM-$\mathcal{L}_{vgg}$ (Reda et al. 2022) 39.79 0.9900 0.0123 0.0128 35.77 0.9782 0.0219 0.0183
FILM-$\mathcal{L}_{s}$ (Reda et al. 2022) 39.68 0.9900 0.0120 0.0124 35.70 0.9781 0.0213 0.0177
LDMVFI (Danier, Zhang, and Bull 2024) 38.68 0.9834 0.0145 0.0130 33.90 0.9703 0.0284 0.0219
PerVFI (Wu et al. 2024) 38.02 0.9831 0.0142 0.0124 34.57 0.9662 0.0245 0.0181
@@
MoMo-10M (Ours) 39.54 0.9896 0.0111 0.0103 35.36 0.9769 0.0204 0.0157
Table A2: Full quantitative results including the fidelity metrics (PSNR, SSIM) on the ‘easy’ and ‘medium’ subsets of SNU-FILM benchmark (Choi et al. 2020) The best results are in bold, and the second best is underlined, respectively.
Method Perception- oriented loss SNU-FILM-hard SNU-FILM-extreme
@@
ABME (Park, Lee, and Kim 2021) 30.58 0.9364 0.0658 0.0496 25.42 0.8639 0.1258 0.0747
XVFIv (Sim, Oh, and Kim 2021) 29.91 0.9073 0.0629 0.0414 24.67 0.8092 0.1257 0.0673
IFRNet-Large (Kong et al. 2022) 30.63 0.9368 0.0562 0.0403 25.27 0.8609 0.1131 0.0638
RIFE (Huang et al. 2022b) 30.10 0.9330 0.0657 0.0443 24.84 0.8534 0.1390 0.0764
FILM-$\mathcal{L}_{1}$ (Reda et al. 2022) 30.42 0.9353 0.0568 0.0441 25.17 0.8593 0.1060 0.0632
AMT-G (Li et al. 2023) 30.70 0.9381 0.0680 0.0506 25.64 0.8658 0.1128 0.0686
EMA-VFI (Zhang et al. 2023) 30.79 0.9386 0.0579 0.0457 25.59 0.8648 0.1099 0.0671
UPRNet-LARGE (Jin et al. 2023) 30.86 0.9377 0.0612 0.0475 25.63 0.8641 0.1109 0.0672
CAIN (Choi et al. 2020) 29.90 0.9292 0.0885 0.0606 24.78 0.8507 0.1790 0.1042
FILM-$\mathcal{L}_{vgg}$ (Reda et al. 2022) 30.34 0.9332 0.0443 0.0282 25.11 0.8557 0.0917 0.0471
FILM-$\mathcal{L}_{s}$ (Reda et al. 2022) 30.29 0.9329 0.0429 0.0268 25.07 0.8550 0.0889 0.0448
LDMVFI (Danier, Zhang, and Bull 2024) 28.51 0.9173 0.0602 0.0379 23.92 0.8372 0.1226 0.0651
PerVFI (Wu et al. 2024) 29.68 0.9287 0.0561 0.0635 25.03 0.8120 0.0902 0.0448
@@
MoMo-10M (Ours) 30.00 0.9308 0.0425 0.0257 24.91 0.8535 0.0882 0.0438
Table A3: Full quantitative results including the fidelity metrics (PSNR, SSIM) on the ‘hard’ and ‘extreme’ subsets of SNU-FILM benchmark (Choi et al. 2020). The best results are in bold, and the second best is underlined, respectively.
Method Perception- oriented loss Middlebury Vimeo90k
@@
ABME (Park, Lee, and Kim 2021) 37.05 0.9845 0.0290 0.0325 36.18 0.9805 0.0213 0.0353
XVFIv (Sim, Oh, and Kim 2021) 36.72 0.9826 0.0169 0.0244 35.07 0.9710 0.0229 0.0354
IFRNet-Large (Kong et al. 2022) 36.27 0.9816 0.0285 0.0366 36.20 0.9808 0.0189 0.0325
RIFE (Huang et al. 2022b) 37.16 0.9853 0.0162 0.0228 35.61 0.9780 0.0223 0.0356
FILM-$\mathcal{L}_{1}$ (Reda et al. 2022) 37.37 0.9838 0.0173 0.0246 35.89 0.9796 0.0197 0.0343
AMT-G (Li et al. 2023) 34.23 0.9708 0.0486 0.0533 36.53 0.9819 0.0195 0.0351
EMA-VFI (Zhang et al. 2023) 38.32 0.9871 0.0151 0.0218 36.45 0.9811 0.0196 0.0343
UPRNet-LARGE (Jin et al. 2023) 38.09 0.9861 0.0150 0.0209 36.42 0.9815 0.0201 0.0342
CAIN (Choi et al. 2020) 35.11 0.9761 0.0254 0.0383 34.65 0.9729 0.0306 0.0483
FILM-${\mathcal{L}_{vgg}}$ (Reda et al. 2022) 37.28 0.9843 0.0096 0.0148 35.62 0.9784 0.0137 0.0229
FILM-$\mathcal{L}_{s}$ (Reda et al. 2022) 37.38 0.9844 0.0093 0.0140 35.71 0.9787 0.0131 0.0224
LDMVFI (Danier, Zhang, and Bull 2024) 34.03 0.9648 0.0195 0.0261 33.09 0.9558 0.0233 0.0327
PerVFI (Wu et al. 2024) 35.00 0.9751 0.0142 0.0163 34.00 0.9675 0.0179 0.0248
MoMo (Ours) 36.77 0.9806 0.0094 0.0126 34.94 0.9756 0.0136 0.0203
MoMo-10M (Ours) 36.52 0.9801 0.0100 0.0139 34.82 0.9752 0.0138 0.0206
Table A4: Full quantitative results including the fidelity metrics (PSNR, SSIM) on Middlebury (Baker et al. 2011) and Vimeo90k (Xue et al. 2019) benchmarks. The best results are in bold, and the second best is underlined, respectively.
Method Perception- oriented loss Xiph-2K Xiph-4K
PSNR SSIM LPIPS DISTS PSNR SSIM LPIPS DISTS
ABME (Park, Lee, and Kim 2021) 36.50 0.9668 0.1071 0.0581 33.72 0.9452 0.2361 0.1108
XVFIv (Sim, Oh, and Kim 2021) 35.17 0.9625 0.0844 0.0418 32.45 0.9274 0.1835 0.0779
IFRNet-Large (Kong et al. 2022) 36.40 0.9646 0.0681 0.0372 33.71 0.9425 0.1364 0.0665
RIFE (Huang et al. 2022b) 36.06 0.9642 0.0918 0.0481 33.21 0.9413 0.2072 0.0915
FILM-$\mathcal{L}_{1}$ (Reda et al. 2022) 36.53 0.9663 0.0906 0.0510 33.83 0.9439 0.1841 0.0884
AMT-G (Li et al. 2023) 36.29 0.9647 0.1061 0.0563 34.55 0.9472 0.2054 0.1005
EMA-VFI (Zhang et al. 2023) 36.74 0.9675 0.1024 0.0550 34.55 0.9486 0.2258 0.1049
UPRNet-LARGE (Jin et al. 2023) 37.13 0.9691 0.1010 0.0553 34.57 0.9388 0.2150 0.1017
CAIN (Choi et al. 2020) 35.18 0.9625 0.1025 0.0533 32.55 0.9398 0.2229 0.0980
FILM-${\mathcal{L}_{vgg}}$ (Reda et al. 2022) 36.29 0.9626 0.0355 0.0238 33.44 0.9356 0.0754 0.0406
FILM-$\mathcal{L}_{s}$ (Reda et al. 2022) 36.30 0.9616 0.0330 0.0237 33.37 0.9323 0.0703 0.0385
LDMVFI (Danier, Zhang, and Bull 2024) 33.82 0.9494 0.0420 0.0163 31.39 0.9214 0.0859 0.0359
PerVFI (Wu et al. 2024) 34.69 0.9541 0.0381 0.0153 32.30 0.9149 0.0858 0.0331
@@
MoMo-10M (Ours) 35.23 0.9548 0.0303 0.0120 32.97 0.9281 0.0638 0.0275
Table A5: Full quantitative results including the fidelity metrics (PSNR, SSIM) on Xiph-2K and Xiph-4K (Montgomery and Lars 1994; Niklaus and Liu 2020). The best results are in bold, and the second best is underlined, respectively.